# WENDY: Gene Regulatory Network Inference with Covariance Dynamics

### v1

@@ 3.3 Optimization formulation for covariance dynamics
Assume we measure the expression levels of $n$ genes for $m$ single cells. When $m<n$, which is common in reality, if we directly calculate the covariance matrix, it will always be degenerate (non-invertible). Therefore, we need to apply a specific method, called graphical lasso, that estimates the covariance matrix $K$ in this case, where $K$ is invertible, and the inverse of $K$ is sparse [26].
For a $1\times n$-dimensional random vector $N=[N_{1},\ldots,N_{n}]$ with invertible covariance matrix $K$, there is a result that $K^{-1}_{i,j}=0$ if and only if the partial Pearson correlation coefficient satisfies [88]
@@
By assuming the expression levels of $n$ genes satisfy a multivariate normal distribution, there is a GRN inference method [60]: gene $i$ and gene $j$ have a direct regulatory relation (direction unknown) if and only if $K^{-1}_{i,j}\neq 0$. Nevertheless, this method assumes that the gene expression is in stationary.
@@
We use the BFGS algorithm to minimize (10). The overall algorithm is presented in Algorithm 1. For the WENDY method, step 4 of Algorithm 1 is much faster than step 2, since the solver terminates after a constant number of iterations. Graphical lasso has time complexity $\mathcal{O}(n^{3})$ [26], which does not depend on the cell number $m$. Therefore, the overall time complexity of WENDY method is $\mathcal{O}(n^{3})$. In Section 4, we will see that in practice, the time cost of WENDY increases with $n$, but not $m$.
@@ 3.4 Theoretical comparison with other methods
(I, II) For an expression level data set with $n$ genes at $T$ time points, SINCERITIES only extracts $n(T-1)$ values, and then applies linear regression. If we transform the data into Scenario 8, we can only obtain $nT$ values before fitting to an ODE system. If we abandon data at other time points to switch to Scenario 2, we lose the temporal information. In comparison, WENDY extracts $n^{2}T/2+nT/2$ values from such data before proceeding to the next step. Therefore, WENDY can extract more information from data. SINCERITIES requires at least $T=3$ time points to work, and requires at least $T=6$ time points to perform well. Methods for Scenario 8 need at least $T=3$ time points to work, and also more time points to work well. In comparison, WENDY only needs $T=2$ time points. This can be explained by information theory: the GRN (without autoregulation) has $n^{2}-n$ independent values. WENDY uses data at $T=2$ time points to extract $n^{2}+n$ values; SINCERITIES needs data at $T=n$ time points to extract $n^{2}-n$ values; methods for Scenario 8 need data at $T=n-1$ time points to extract $n^{2}-n$ values.
@@
(IV) Harissa and CARDAMOM [35, 77] have the same idea: for Scenario 2 data, build a nonlinear regression model under complicated pre-processing. Given the single-cell expression data $D_{t}$ at time $t$, they can calculate the regression error $R(D_{t},A_{t})$ for any possible GRN $A_{t}$. Different from (III), they do not directly minimize $R(D_{t},A_{t})$, but add a calibration term that prevent the current inferred GRN $A_{t}$ from being too different with the inferred GRN $A_{t-1}$ for the previous time point $t-1$:
@@ 4 Performance evaluation on synthetic data
In this section, we use two synthetic data sets with different settings to test the performance of WENDY and four other GRN inference methods that work for different data types: GENIE3, dynGENIE3, NonlinearODEs, SINCERITIES. WENDY ranks the second for the DREAM4 data set, and ranks the first for the SINC data set.
@@ 4.1 Methods and measurements
GENIE3 [38] works for Scenario 2, single-cell level one-time expression data at stationary. The idea is to infer the level of one gene by the levels of other genes using random forest or extra trees. All edges can be inferred, but sometimes the direction is hard to determine.
dynGENIE3 [37] is the revised version of GENIE3 that works for time series data, such as Scenario 5 (and 10), single-cell level time series expression data at stationary or after general interventions. The idea is basically the same as GENIE3, but here it infers the level of one gene at a later time by the levels of other genes at an earlier time. Therefore, all edges including the directions can be determined. dynGENIE3 requires the correspondence of cells at different time points. This means that it cannot be applied to Scenario 9 data directly. Therefore, dynGENIE3 is included just for completeness, not as a main comparison target.
NonlinearODEs [51] works for Scenario 8, bulk level time series expression data measured after general interventions. The idea is to fit the data to a nonlinear ODE model. It uses XGBoost to determine the importance score of each edge. There is a tunable parameter $\alpha$ that can be chosen manually or automatically, and we use the from_data mode to determine the parameter $\alpha$ automatically. All edges including the directions can be determined.
SINCERITIES [61] works for Scenario 9, single-cell level time series expression data measured after general interventions, where the joint distribution at different time points is unknown. The idea is to calculate the distance between the distributions of the same gene at two time points, and then applies linear regression. All edges including the directions can be determined.
@@
To test a GRN inference method under different circumstances, a common practice is to use synthetic data [61, 51, 37]. The data are generally generated by numerically simulating an SDE system. We will test different inference methods on two data sets: DREAM4 and SINC.
Each inference method obtains a calculated GRN matrix $A^{\prime}$, whose entries take values in $\mathbb{R}$. To generate the synthetic data, there is a true GRN matrix $A$, which can take values $/-1$ 0 / - 1. To evaluate the inference result $A^{\prime}$ with the true GRN matrix $A$, we calculate two measurements, AUROC and AUPR [61]. AUROC is the area under the curve of true positive rate versus false positive rate, and AUPR is the area under the curve of precision versus recall. They evaluate how the inferred GRN fits the true GRN from different perspectives. Since $A^{\prime}_{i,i}\neq 0$ does not necessarily mean autoregulation, we do not compare the diagonal elements of $A$ and $A^{\prime}$. See Algorithm 2 for the calculation procedure of such quantities.
@@
    Initiation: true positive (TP) is 00; true negative (TN) is the number of 00 in all off-diagonal elements of $A$; false positive (FP) is 00; false negative (FN) is the number of $/-1$ 1 in all off-diagonal elements of $A$; TPR0=FPR0=Precision0=Recall0=0
@@
    Sort the off-diagonal elements of $A^{\prime}$ by its abstract value in descending order as an array, denoted as Val
@@
    For index $k=1,\ldots,n^{2}-n$:
    $A^{\prime}_{i,j}=$ Valk
    If ($A^{\prime}_{i,j}\geq 0$ and $A_{i,j}=1$) or ($A^{\prime}_{i,j}<0$ and $A_{i,j}=-1$):
    TP=TP+1; FN=FN-1
    Elif ($A^{\prime}_{i,j}<0$ and $A_{i,j}=1$) or ($A^{\prime}_{i,j}\geq 0$ and $A_{i,j}=-1$):
    FP=FP+1; FN=FN-1
    Else ($A_{i,j}=0$):
    FP=FP+1; TN=TN-1
    End of if
    Calculate TPRk=TP/(TP+FN); FPRk=FP/(FP+TN); Precisionk=TP/(TP+FP); Recallk=TP/(TP+FN)=TPRk
@@
    Calculate AUROC, the area under the curve of TPR vs. FPR; calculate AUPR, the area under the curve of Precision vs. Recall
  6. 6.
@@ 4.2 DREAM4 data set
The most common synthetic data set used to evaluate GRN inference methods is DREAM4 – In Silico Network Challenge [52]. It has multiple challenges, each with multiple data types. We will consider the two in silico challenges and use only the time series data set. There are $5$ GRNs with $10$ genes, each accompanied by $5$ stochastic trajectories at $21$ time points. There are also $5$ GRNs with $100$ genes, each accompanied by $10$ stochastic trajectories at $21$ time points. The time points are equally distributed: 0, 50, 100, 150, 200, 250, 300, 350, 400, 450, 500, 550, 600, 650, 700, 750, 800, 850, 900, 950, 1000. Each GRN is represented by a matrix $A$, where $A_{i,j}=1$ or $A_{i,j}=0$, which means that gene $i$ can or cannot regulate gene $j$. Also, $A_{i,i}=0$ for each gene $i$. These two data sets are denoted as DREAM4 (10 genes) and DREAM4 (100 genes). DREAM4 data are generated by GeneNetWeaver software [69], which integrates some unknown SDE system.
@@
(2) For SINCERITIES, we regard DREAM4 data as Scenario 9 data by ignoring the cell correspondence between different time points. For DREAM4 data (10 genes), we use the data at $6$ time points: $t=$(100, 150, 200, 250, 300, 350). For DREAM4 data (100 genes), we use the data at $5$ time points: $t=$(750, 800, 850, 900, 950).
@@
(4) For GENIE3, we regard DREAM4 data as Scenario 2 data by considering only one time point. For DREAM4 data (10 genes), we use the data at $t=$350. For DREAM4 data (100 genes), we use the data at $t=$450.
(5) For dynGENIE3, we regard DREAM4 data as Scenario 10 data. For DREAM4 data (10 genes), we use the data at $13$ time points: $t=$(0, 50, 100, 150, 200, 250, 300, 350, 400, 450, 500, 550, 600). For DREAM4 data (100 genes), we use the data at $18$ time points: $t=$(0, 50, 100, 150, 200, 250, 300, 350, 400, 450, 500, 550, 600, 650, 700, 750, 800, 850).
See Table 2 for the results, where the AUROC and AUPR are averaged over all 5 GRNs in the same data set. We can see that on average, WENDY is slightly better than NonlinearODEs and GENIE3, which are slightly better than SINCERITIES, while dynGENIE3 is significantly better than all other methods. This is not surprising, since dynGENIE3 can utilize the cell correspondence (joint distribution) of different time points, which is not realistic, but contains more information.
Besides, some results are lower than those values reported in corresponding papers [38, 37, 51], since we only use the time series data, not combining with other data types, and we do not manually fine-tune the parameters accordingly.
@@
DREAM4 AUROC 0.64 0.61 0.64 0.63 0.76
(10 genes) AUPR 0.29 0.21 0.28 0.24 0.46
DREAM4 AUROC 0.62 0.55 0.58 0.64 0.74
(100 genes) AUPR 0.05 0.03 0.03 0.04 0.14
Total 1.60 1.40 1.53 1.55 2.10
@@ 4.3 SINC data set
Pinna et al. and Papili Gao et al. [64, 61] use the following SDE to generate synthetic data:
@@
Here $X_{i}(t)$ is the expression level of gene $i$ at time $t$, and $W_{j}(t)$ is a standard Brownian motion, independent with other $W_{i}(\tau)$. $A_{i,j}$ describes the GRN. $V=30$, $\beta=1$, $\theta=0.2$, $\sigma=0.1$. When $t\to\infty$, each $X_{i}(t)$ will converge to the stationary value and fluctuate slightly around it [93, 14, 15].
@@
For Eq. 11, we simulate the system with the Euler-Maruyama method [42] for time $t\in[0,1]$ with time step $\Delta t=0.01$. This means treating $\mathrm{d}X_{j}(t)$ as $X_{j}(t+\Delta t)-X_{j}(t)$, $\mathrm{d}t$ as $\Delta t$, and $\mathrm{d}W_{j}(t)$ as a normal random variable $\mathcal{N}(0,\Delta t)$.
For the GRN matrix $A$, we use the $40$ networks in Papili Gao et al. [61]: $10$ Escherichia coli networks with $10$ genes, $10$ E. coli networks with $20$ genes, $10$ Saccharomyces cerevisiae (yeast) networks with $10$ genes, $10$ yeast networks with $20$ genes. Each network has $A_{i,j}=1/0/-1$, and $A_{i,i}=0$ for each $i$.
@@
To test the performance of different methods on SINC data, we compare different settings of the same method and choose the best one. Specifically, we find that for methods that can use data from multiple time points, it is not always good to use data from all $11$ time points. Instead, data from some time points should be abandoned. Here we list the best settings for each method on each data set:
(1) For WENDY, we regard SINC data as Scenario 9 data. For SINC data (10 genes), we use the data at $t=0.0$ and $t=3.0$ without the cell correspondence between different time points. For SINC data (20 genes), we use the data at $t=2.7$ and $t=3.0$.
(2) For SINCERITIES, we regard SINC data as Scenario 9 data by ignoring the cell correspondence between different time points. For SINC data (10 genes), we use the data at $8$ time points: $t=$(0.9, 1.2, 1.5, 1.8, 2.1, 2.4, 2.7, 3.0). For SINC data (100 genes), we use the data at $9$ time points: $t=$(0.6, 0.9, 1.2, 1.5, 1.8, 2.1, 2.4, 2.7, 3.0). This fits with Papili Gao et al.’s paper, as they also abandon the data for $t\leq 0.5$.
(3) For NonlinearODEs, we regard SINC data as Scenario 8 data by taking average of gene expression levels over different cells. For SINC data (10 genes), we use the data at $10$ time points: $t=$(0.3, 0.6, 0.9, 1.2, 1.5, 1.8, 2.1, 2.4, 2.7, 3.0). For SINC data (100 genes), we use the data at $8$ time points: $t=$(0.3, 0.6, 0.9, 1.2, 1.5, 1.8, 2.1, 2.4).
(4) For GENIE3, we regard SINC data as Scenario 2 data by considering only one time point. For both SINC data (10 genes) and SINC data (100 genes), we use the data at $t=0.3$.
(5) For dynGENIE3, we regard SINC data as Scenario 10 data. For both SINC data (10 genes) and SINC data (100 genes), we use the data at $10$ time points: $t=$(0.0, 0.3, 0.6, 0.9, 1.2, 1.5, 1.8, 2.1, 2.4, 2.7).
We apply WENDY, SINCERITIES, NonlinearODEs, GENIE3, and dynGENIE3 to SINC data sets and compare the corresponding AUROC and AUPR. Each AUROC or AUPR value is averaged over 2000 simulations (20 GRNs, each with 100 groups of data). See Table 3 for the results. We can see that WENDY is significantly better than all other methods. Here dynGENIE3 has the worst performance, possibly because it cannot determine the sign (activation/inhibition) of a regulation.
For different values of cell/trajectory number $m$, we can see that the performance metrics generally increase with $m$, meaning that more cells provide more information.
@@
SINC $m=10$ AUROC 0.59 0.44 0.47 0.46 0.43
(10 genes) AUPR 0.14 0.11 0.11 0.11 0.10
SINC $m=10$ AUROC 0.76 0.42 0.51 0.52 0.40
(20 genes) AUPR 0.16 0.05 0.08 0.08 0.04
SINC $m=30$ AUROC 0.67 0.46 0.49 0.50 0.44
(10 genes) AUPR 0.19 0.12 0.12 0.15 0.10
SINC $m=30$ AUROC 0.81 0.47 0.53 0.47 0.40
(20 genes) AUPR 0.21 0.07 0.08 0.07 0.04
SINC $m=100$ AUROC 0.69 0.48 0.50 0.55 0.42
(10 genes) AUPR 0.22 0.13 0.12 0.19 0.10
SINC $m=100$ AUROC 0.82 0.48 0.54 0.54 0.39
(20 genes) AUPR 0.23 0.08 0.08 0.09 0.05
Total 5.49 3.31 3.63 3.73 2.91
Notice that WENDY is not always better in all settings. For example, if we only have the data at $t=0.0$ and $t=0.1$ for SINC data with $m=100$, then WENDY has AUROC 0.77 and AUPR 0.34 for SINC ($10$ genes) data, and AUROC 0.83 and AUPR 0.29 for SINC ($20$ genes) data, while GENIE3 has AUROC 0.77 and AUPR 0.32 for SINC ($10$ genes) data, and AUROC 0.92 and AUPR 0.50 for SINC ($20$ genes) data. Besides, for the setting in Table 3, if we reduce $m$ to $5$, then the performance of WENDY is also no better than GENIE3.

### v2

@@ 3.3 Optimization formulation for covariance dynamics
Assume we measure the expression levels of $n$ genes for $m$ single cells. When $m<n$, which is common in reality, if we directly calculate the covariance matrix, it will always be degenerate (non-invertible). Therefore, we need to apply a specific method, called graphical lasso, that estimates the covariance matrix $K$ in this case, where $K$ is invertible, and the inverse of $K$ is sparse (Friedman et al.,, 2008).
For a $1\times n$-dimensional random vector $N=[N_{1},\ldots,N_{n}]$ with invertible covariance matrix $K$, there is a result that $K^{-1}_{i,j}=0$ if and only if the partial Pearson correlation coefficient satisfies (Wasserman,, 2004)
@@
By assuming the expression levels of $n$ genes satisfy a multivariate normal distribution, there is a GRN inference method (Nouri et al.,, 2023): gene $i$ and gene $j$ have a direct regulatory relation (direction unknown) if and only if $K^{-1}_{i,j}\neq 0$. Nevertheless, this method assumes that the gene expression is in stationary.
@@
We use the BFGS algorithm to minimize (10). The overall algorithm is presented in Algorithm 1. For the WENDY method, step 4 of Algorithm 1 is much faster than step 2, since the solver terminates after a constant number of iterations. Graphical lasso has time complexity $\mathcal{O}(n^{3})$ (Friedman et al.,, 2008), which does not depend on the cell number $m$. Therefore, the overall time complexity of WENDY method is $\mathcal{O}(n^{3})$. In Section 4, we will see that in practice, the time cost of WENDY increases with $n$, but not $m$.
@@ 3.4 Theoretical comparison with other methods
(I, II) For an expression level data set with $n$ genes at $T$ time points, SINCERITIES only extracts $n(T-1)$ values, and then applies linear regression. If we transform the data into Scenario 8, we can only obtain $nT$ values before fitting to an ODE system. If we abandon data at other time points to switch to Scenario 2, we lose the temporal information. In comparison, WENDY extracts $n^{2}T/2+nT/2$ values from such data before proceeding to the next step. Therefore, WENDY can extract more information from data. SINCERITIES requires at least $T=4$ time points to work normally (with $3$ time points, SINCERITIES always produces an all-zero matrix), and requires at least $T=6$ time points to perform well. Methods for Scenario 8 need at least $T=3$ time points to work, and also more time points to work well. In comparison, WENDY only needs $T=2$ time points. This can be explained by information theory: the GRN (without autoregulation) has $n^{2}-n$ independent values. WENDY uses data at $T=2$ time points to extract $n^{2}+n$ values; SINCERITIES needs data at $T=n$ time points to extract $n^{2}-n$ values; methods for Scenario 8 need data at $T=n-1$ time points to extract $n^{2}-n$ values.
@@
(IV) Harissa and CARDAMOM (Herbach,, 2023; Ventre,, 2021) have the same idea: for Scenario 2 data, build a nonlinear regression model under complicated pre-processing. Given the single-cell expression data $D_{t}$ at time $t$, they can calculate the regression error $R(D_{t},A_{t})$ for any possible GRN $A_{t}$. Different from (III), they do not directly minimize $R(D_{t},A_{t})$, but add a calibration term that prevent the current inferred GRN $A_{t}$ from being too different with the inferred GRN $A_{t-1}$ for the previous time point $t-1$:
@@ 4 Performance evaluation on synthetic data
In this section, we use two synthetic data sets with different settings to test the performance of WENDY and four other GRN inference methods that work for different data types: GENIE3, dynGENIE3, NonlinearODEs, SINCERITIES. WENDY ranks second for the DREAM4 data set, and ranks third for the SINC data set.
@@ 4.1 Methods and measurements
GENIE3 (Huynh-Thu et al.,, 2010) works for Scenario 2, single-cell level one-time expression data at stationary. The idea is to infer the level of one gene by the levels of other genes using random forest or extra trees. All edges can be inferred, but sometimes the direction is hard to determine.
dynGENIE3 (Huynh-Thu and Geurts,, 2018) is the revised version of GENIE3 that works for time series data, such as Scenario 5 (and 10), single-cell level time series expression data at stationary or after general interventions. The idea is basically the same as GENIE3, but here it infers the level of one gene at a later time by the levels of other genes at an earlier time. Therefore, all edges including the directions can be determined. dynGENIE3 requires the correspondence of cells at different time points. This means that it cannot be applied to Scenario 9 data directly. Therefore, dynGENIE3 is included just for completeness, not as a main comparison target.
NonlinearODEs (Ma et al.,, 2020) works for Scenario 8, bulk level time series expression data measured after general interventions. The idea is to fit the data to a nonlinear ODE model. It uses XGBoost to determine the importance score of each edge. There is a tunable parameter $\alpha$ that can be chosen manually or automatically, and we use the from_data mode to determine the parameter $\alpha$ automatically. All edges including the directions can be determined.
SINCERITIES (Papili Gao et al.,, 2018) works for Scenario 9, single-cell level time series expression data measured after general interventions, where the joint distribution at different time points is unknown. The idea is to calculate the distance between the distributions of the same gene at two time points, and then applies linear regression. All edges including the directions can be determined.
@@
To test a GRN inference method under different circumstances, a common practice is to use synthetic data (Papili Gao et al.,, 2018; Ma et al.,, 2020; Huynh-Thu and Geurts,, 2018). The data are generally generated by numerically simulating an SDE system. We will test different inference methods on two data sets: DREAM4 and SINC.
Each inference method obtains a calculated GRN matrix $A^{\prime}$, whose entries take values in $\mathbb{R}$. To generate the synthetic data, there is a true GRN matrix $A$, which can take values $/-1$ 0 / - 1. To evaluate the inference result $A^{\prime}$ with the true GRN matrix $A$, we calculate two measurements, AUROC and AUPR (Papili Gao et al.,, 2018). AUROC is the area under the curve of true positive rate versus false positive rate, and AUPR is the area under the curve of precision versus recall. They evaluate how the inferred GRN fits the true GRN from different perspectives. Since $A^{\prime}_{i,i}\neq 0$ does not necessarily mean autoregulation, we do not compare the diagonal elements of $A$ and $A^{\prime}$.
AUROC and AUPR are originally defined for binary classification problems, where the ground truth can only take two values (e.g., 1 and 0), but the true GRN in our problem can take three values. We need to generalize the definitions of AUROC and AUPR. AUROC has another equivalent definition: for all data point pairs $(i,j)$ and $(p,q)$ with $A_{i,j}\neq A_{p,q}$, AUROC is the proportion of data point pairs that $(A_{i,j}-A^{\prime}_{i,j})(A_{p,q}-A^{\prime}_{p,q})>0$. This definition works for our ternary case. Notice that this ternary definition of AUROC is equivalent to calculate the binary AUROC for data points with $A_{i,j}=1$ or $A_{p,q}=0$, and weight it by the number of data point pairs $A_{i,j}=1,A_{p,q}=0$, and repeat this to $A_{i,j}=1,A_{p,q}=-1$ and $A_{i,j}=0,A_{p,q}=-1$. Then AUPR can be defined similarly. See Algorithm 2 for the calculation procedure of such quantities. When the true GRN can only take two values, these ternary definitions degenerate back to the original binary AUROC and AUPR. Besides, these ternary definitions also lead to values between 0 and 1, where 1 means perfect ordering.
@@
    Setup:
    $\mathcal{U}=\{(i,j)\mid A_{i,j}=1,i\neq j\}$, $u=|\mathcal{U}|$
    $\mathcal{Z}=\{(i,j)\mid A_{i,j}=0,i\neq j\}$, $z=|\mathcal{Z}|$
    $\mathcal{D}=\{(i,j)\mid A_{i,j}=-1,i\neq j\}$, $d=|\mathcal{D}|$
@@
    Calculate
    $\mathrm{AUROC}_{1/0}^{\mathrm{binary}}$ and $\mathrm{AUPR}_{1/0}^{\mathrm{binary}}$ for data points in $\mathcal{U}$ and $\mathcal{Z}$
    $\mathrm{AUROC}_{1/-0}^{\mathrm{binary}}$ and $\mathrm{AUPR}_{1/-0}^{\mathrm{binary}}$ for data points in $\mathcal{U}$ and $\mathcal{D}$
    $\mathrm{AUROC}_{0/-1}^{\mathrm{binary}}$ and $\mathrm{AUPR}_{0/-1}^{\mathrm{binary}}$ for data points in $\mathcal{Z}$ and $\mathcal{D}$
@@
    Average
    $\mathrm{AUROC}^{\mathrm{ternary}}=\frac{uz\ \mathrm{AUROC}_{1/0}^{\mathrm{binary}}}{uz+ud+zd}+\frac{ud\ \mathrm{AUROC}_{1/-1}^{\mathrm{binary}}}{uz+ud+zd}+\frac{zd\ \mathrm{AUROC}_{0/-1}^{\mathrm{binary}}}{uz+ud+zd}$
    $\mathrm{AUPR}^{\mathrm{ternary}}=\frac{uz\ \mathrm{AUPR}_{1/0}^{\mathrm{binary}}}{uz+ud+zd}+\frac{ud\ \mathrm{AUPR}_{1/-1}^{\mathrm{binary}}}{uz+ud+zd}+\frac{zd\ \mathrm{AUPR}_{0/-1}^{\mathrm{binary}}}{uz+ud+zd}$
@@ 4.2 DREAM4 data set
The most common synthetic data set used to evaluate GRN inference methods is DREAM4 – In Silico Network Challenge (Marbach et al.,, 2012). It has multiple challenges, each with multiple data types. We will consider the two in silico challenges and use only the time series data set. There are $5$ GRNs with $10$ genes, each accompanied by $5$ stochastic trajectories at $21$ time points. There are also $5$ GRNs with $100$ genes, each accompanied by $10$ stochastic trajectories at $21$ time points. The time points are equally distributed: 0, 50, 100, 150, 200, 250, 300, 350, 400, 450, 500, 550, 600, 650, 700, 750, 800, 850, 900, 950, 1000. Each GRN is represented by a matrix $A$, where $A_{i,j}=1$ or $A_{i,j}=0$, which means that gene $i$ can or cannot regulate gene $j$. Also, $A_{i,i}=0$ for each gene $i$. These two data sets are denoted as DREAM4 (10 genes) and DREAM4 (100 genes). DREAM4 data are generated by GeneNetWeaver software (Schaffter et al.,, 2011), which integrates some unknown SDE system.
@@
(2) For SINCERITIES, we regard DREAM4 data as Scenario 9 data by ignoring the cell correspondence between different time points. For DREAM4 data (10 genes), we use the data at $8$ time points: $t=$(0, 50, 100, 150, 200, 250, 300, 350). For DREAM4 data (100 genes), we use the data at $5$ time points: $t=$(750, 800, 850, 900, 950).
@@
(4) For GENIE3, we regard DREAM4 data as Scenario 2 data by considering only one time point. For DREAM4 data (10 genes), we use the data at $t=$450. For DREAM4 data (100 genes), we use the data at $t=$350.
(5) For dynGENIE3, we regard DREAM4 data as Scenario 10 data. For DREAM4 data (10 genes), we use the data at $18$ time points: $t=$(0, 50, 100, 150, 200, 250, 300, 350, 400, 450, 500, 550, 600, 650, 700, 750, 800, 850). For DREAM4 data (100 genes), we use the data at $17$ time points: $t=$(0, 50, 100, 150, 200, 250, 300, 350, 400, 450, 500, 550, 600, 650, 700, 750, 800).
See Table 2 for the results, where the AUROC and AUPR are averaged over all 5 GRNs in the same data set. We can see that on average, WENDY is slightly better than NonlinearODEs and GENIE3, which are significantly better than SINCERITIES, while dynGENIE3 is significantly better than all other methods. This is not surprising, since dynGENIE3 can utilize the cell correspondence (joint distribution) of different time points, which is not realistic, but contains more information.
Besides, some results are lower than those values reported in corresponding papers (Huynh-Thu et al.,, 2010; Huynh-Thu and Geurts,, 2018; Ma et al.,, 2020), since we only use the time series data, not combining with other data types, and we do not manually fine-tune the parameters accordingly.
@@
DREAM4 AUROC 0.65 0.52 0.64 0.64 0.75
(10 genes) AUPR 0.34 0.19 0.32 0.27 0.54
DREAM4 AUROC 0.59 0.55 0.57 0.64 0.74
(100 genes) AUPR 0.05 0.03 0.04 0.05 0.14
Total 1.63 1.29 1.57 1.60 2.17
@@ 4.3 SINC data set
Pinna et al., (2010) and Papili Gao et al., (2018) use the following SDE to generate synthetic data:
@@
Here $X_{i}(t)$ is the expression level of gene $i$ at time $t$, and $W_{j}(t)$ is a standard Brownian motion, independent with other $W_{i}(\tau)$. $A_{i,j}$ describes the GRN. $V=30$, $\beta=1$, $\theta=0.2$, $\sigma=0.1$. When $t\to\infty$, each $X_{i}(t)$ will converge to the stationary value and fluctuate slightly around it (Yang and Cheng,, 2021; Cheng and Qian,, 2021; Cheng et al.,, 2021).
@@
For Eq. 11, we simulate the system with the Euler-Maruyama method (Kloeden et al.,, 1992) for time $t\in[0,1]$ with time step $\Delta t=0.01$. This means treating $\mathrm{d}X_{j}(t)$ as $X_{j}(t+\Delta t)-X_{j}(t)$, $\mathrm{d}t$ as $\Delta t$, and $\mathrm{d}W_{j}(t)$ as a normal random variable $\mathcal{N}(0,\Delta t)$.
For the GRN matrix $A$, we use the $40$ networks in Papili Gao et al., (2018): $10$ Escherichia coli networks with $10$ genes, $10$ E. coli networks with $20$ genes, $10$ Saccharomyces cerevisiae (yeast) networks with $10$ genes, $10$ yeast networks with $20$ genes. Each network has $A_{i,j}=1/0/-1$, and $A_{i,i}=0$ for each $i$.
@@
To test the performance of different methods on SINC data, we compare different settings of the same method and choose the best one. For WENDY, we fix the first time point to be 0.0. For SINCERITIES, NonlinearODEs, and dynGENIE3, we consider contiguous time points that start from 0.0. Specifically, we find that for methods that can use data from multiple time points, it is not always good to use data from all $11$ time points. Instead, data from some time points should be abandoned. Here we list the best settings for each method on each data set:
(1) For WENDY, we regard SINC data as Scenario 9 data. For both SINC data (10 genes)and SINC data (20 genes), we use the data at $t=0.0$ and $t=0.3$.
(2) For SINCERITIES, we regard SINC data as Scenario 9 data by ignoring the cell correspondence between different time points. For both SINC data (10 genes) and SINC data (20 genes), we use the data at $4$ time points: $t=$(0.0, 0.3, 0.6, 0.9).
(3) For NonlinearODEs, we regard SINC data as Scenario 8 data by taking average of gene expression levels over different cells. For both SINC data (10 genes) and SINC data (20 genes), we use the data at $10$ time points: $t=$(0.0, 0.3, 0.6, 0.9, 1.2, 1.5, 1.8, 2.1, 2.4, 2.7).
(4) For GENIE3, we regard SINC data as Scenario 2 data by considering only one time point. For both SINC data (10 genes) and SINC data (20 genes), we use the data at $t=0.3$.
(5) For dynGENIE3, we regard SINC data as Scenario 10 data. For SINC data (10 genes), we use the data at $4$ time points: $t=$(0.0, 0.3, 0.6, 0.9). For SINC data (20 genes), we use the data at $3$ time points: $t=$(0.0, 0.3, 0.6).
We apply WENDY, SINCERITIES, NonlinearODEs, GENIE3, and dynGENIE3 to SINC data sets and compare the corresponding AUROC and AUPR. Each AUROC or AUPR value is averaged over 2000 simulations (20 GRNs, each with 100 groups of data). See Table 3 for the results. We can see that WENDY is better than SINCERITIES and NonlinearODEs, but weaker than GENIE3 and dynGENIE3 (which does not work for Scenario 9 data).
For different values of cell/trajectory number $m$, we can see that most performance metrics increase with $m$, meaning that more cells provide more information.
@@
SINC $m=10$ AUROC 0.51 0.53 0.49 0.60 0.56
(10 genes) AUPR 0.33 0.32 0.34 0.38 0.36
SINC $m=10$ AUROC 0.50 0.50 0.49 0.54 0.51
(20 genes) AUPR 0.09 0.07 0.09 0.10 0.09
SINC $m=30$ AUROC 0.52 0.53 0.49 0.64 0.59
(10 genes) AUPR 0.33 0.32 0.34 0.40 0.37
SINC $m=30$ AUROC 0.50 0.50 0.49 0.58 0.53
(20 genes) AUPR 0.08 0.08 0.09 0.12 0.09
SINC $m=100$ AUROC 0.58 0.49 0.50 0.70 0.61
(10 genes) AUPR 0.35 0.32 0.34 0.44 0.37
SINC $m=100$ AUROC 0.51 0.47 0.49 0.64 0.55
(20 genes) AUPR 0.09 0.08 0.09 0.14 0.10
Total 4.39 4.21 4.24 5.28 4.73